Agenda/Minutes of the VIS Meeting on 2016/8/5
2016/8/5 13:30 - 14:30 @ Room 210
Participants:Fabian, Mark, Ishizaki, Tanaka, Ohishi, Sato, Okutomi, Hirata, Aso, Akutsu
Simulation: Calculating horizontal-yaw couplings from the asymmetry. Send the results by 8/10 (Okutomi).
- EQ stop: The date for the announce of bidding is the 17th. The specification document and drawings should be fixed by the 10th.
- Installation procedure: The document is under preparation. The review will be held on the 10th.
- LVDT test bench was ordered.
- Dummy payload design is done. Ishizaki-san is making 2D drawings. To be ordered with the EQ stops.
- Drawings for the release sensor were finished. To be ordered with the EQ stops.
- Dummy BS gluing will be done in the next week
- Type-B team will start working in Kamioka from 8/15
- Zhao-kun will help us from 8/15
- Will ask Yokokawa-kun (Toyama University) for help from 8/22 - 9/16 (Aso)
- Will discuss with Liu-kun about the possibility of doing Type-B simulation at the F2F meeting
- The real BS mirror is ready. We can do the gluing work whenever we want.
Type B Past Week Summary
- Akutsu-san provided a set of cleanroom gear (shoes only slightly too tight!) for use by Mark this week. (And Sakamoto-san sent two more sets from Kamioka to Kashiwa!)
- Fabian did a fit test of the winches onto the IM.
- The RM, IM and IRM were weighed, and wrapped/packed for shipment to Kamioka.
- Mark made a checklist of all the parts of the BS mirror box and gluing jig, to help with collecting the parts from Kashiwa.
- Mark, Fabian and Hirata-san retrieved the dummy BS and mirror box from Kashiwa. The dummy BS was stripped of accessories and moved to the ATC fume hood. A combination of acetone and a stainless steel brush was effective in removing the residual glue in the first area worked on.
- Fabian designed an extra 600 g of trim masses for the IM and Hirata-san made drawings. Fukushima-san of ATC made some suggestions for improvement and Fabian will add them. The design allows for the masses to be added to the IM even after it has been placed in the assembly frame and the dummy optic and RM have been hung, so it should not delay work.
- Mark worked on the schedule, making the next few weeks more realistic and incorporating the new plan for the SR RMs (the test will be used for SR2 but two more will be made from scratch next financial year).
- Fabian worked on the payload assembly procedure.
- Hirata-san washed 7 OSEMs for Akutsu-san, and various other parts for Yamamoto-san of ICRR.
- Hirata-san removed the remaining glue from the BS primary and secondary prisms, as well as the prisms removed from the PR hanging test mirror. (The small sapphire prisms for the PR were not intended to be reused and one had already broken at Kashiwa, but the metal PR secondaries can and will be used again.)
- Type B ongoing issues
Akutsu-san is working to get wide-mouth OSEMs (2 assembled; another few were assembled and a batch of 7 were washed.
Longer cables for BF->PI and SF->PI.
- Connectors RECEIVED, cables being made.
- Lower breadboard stuff
- Installation to be incorporated into the procedure document. Mark/Takahashi to discuss.
- To-do list during Mark's vacation
- Weigh IM/RM/IRM (Fabian, Hirata) - DONE
- New trim mass - ORDERED.
- Pack/ship IM/RM/IRM to Kamioka (Hirata/Fabian) - READY TO SHIP.
- Analyze geophone data (Fabian/Shoda) - DONE.
- Remove glue from dummy BS face (Fabian) - MOSTLY DONE.
- Reglue flags and prisms (Tatsumi, Fabian) - from Friday 8/5.
- Clean wire for BS and RM (Hirata) - DONE
- Get new winch parts for new wire diameters - ORDERED, expected 8/5.
- Do fit/function test of winches (Fabian/Hirata) - fit test DONE, function test from 8/8
- Ship winches to Kamioka (Hirata/Fabian)
- Work on BS simulation (Fabian)
- Do Inventor assembly for SR2 in chamber and on assembly frame (Hirata) - ONGOING.
- Write more of payload assembly procedure (Fabian) - ONGOING
- Write more of main assembly procedure (Mark) - ONGOING
- Weigh IM/RM/IRM (Fabian, Hirata) - DONE
- Longer term to-do list
- Do BS Simulink model (Mark, Kokeyama)
- Do 2D drawings of SR2 parts (Hirata)
- Redesign SRx IM wire clamps for larger prisms, remachined RMs
- Order SRM/PRM parts (Hirata/Shoda)
- Order SRx fasteners (Hirata)
- Updated the document tree for Type-Bp
Mirapro sent us the design of the assembly frame -> seems OK
- Asked Mirapro to design the optical bench adapter
- Obuchi-san did the structural analysis of the BF damper part. BFRM has 40Hz lowest resonance. Do simulation to see if this will be a problem or not.
- Simulation: Checked the damping time without local damping
- LVDT check: Working with Sakon-san to test the Virgo LVDT-Actuator
- Started the preparation of work place in the clean room to test the magnet gluing jigs.
Watchdog system (Tanaka)
About the connection and system boundary of the host system.PDF
- No issue
Site work plan for the next week
- Mon: Check existing parts (Shoda)
- Tue: Check existing parts (Shoda)
- Wed: Check existing parts (Shoda)
- Travel (Week of 8/8):
- Shoda: 8-10
- Travel (Week of 8/15):
- Mark: 15-19
- Fabian: 15-19
2016/8/19(Fri) 13:30 -